OpenStudy (anonymous):

I need help. x^2+4x-32=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

factor as \[xZ^2+4x-32=(x-4)(x+8)\] \[(x+4)(x-8)=0\] \[x+4=0\] \[x=-4\] \[x-8=0\] \[x=8\]

OpenStudy (anonymous):

thaks